OpenStudy (anonymous):

find the largest integer that is less than log 10(220,440)

OpenStudy (phi):

You could type log(220440)= in the google search window to find this number or use a calculator or know that \(10^5 \)=100,000 and \(10^6\)=1,000,000

OpenStudy (phi):

knowing that \(10^5\)=100,000 and \(10^6\)=1,000,000 common sense tells you that if \(10^x\)= 220,440, then x must be bigger than 5 but less than 6